OpenStudy (dr0zier99):

mike wants to make meatloaf. his recipe uses a total of 7 pounds of meat

OpenStudy (dr0zier99):

if he uses 3 to 1 ratio of beef to pork how much pork will he use

OpenStudy (dr0zier99):

enter your answer as a mixed number in simplest terms

OpenStudy (anonymous):

i have no idea lol sry

OpenStudy (dr0zier99):

its ok

OpenStudy (dr0zier99):

@TheSmartOne

TheSmartOne (thesmartone):

3x + 1x = 4x 4x = 7 x = 7/4 3(7/4) = ? 1(7/4) = ? Those are your two answers :P

OpenStudy (dr0zier99):

thank you oh smart one

TheSmartOne (thesmartone):

no problem :D